OpenStudy (anonymous):

The operator of a boat wishes to cross a 5km wide river that is flowing to the east at 10 meter per second he wishes to reach the exact point on the opposite shore 15 minutes after starting at what speed and what direction should the boat travel

OpenStudy (anonymous):

L

OpenStudy (irishboy123):



OpenStudy (irishboy123):

 the E-W component of his velocity will have to offset the river's flow: \(v \sin \phi = 10\) the N-S component will have to get him across in 15 minutes: \(v \cos \phi = \dfrac{5000}{15 \times 60}\)